# On the best approximation by finite Gaussian mixtures

### v1

@@ Abstract
We consider the problem of approximating a general Gaussian location mixture by finite mixtures. The minimum order of finite mixtures that achieve a prescribed accuracy (measured by various $f$-divergences) is determined within constant factors for the family of mixing distributions with compactly support or appropriate assumptions on the tail probability including subgaussian and subexponential. While the upper bound is achieved using the technique of local moment matching, the lower bound is established by relating the best approximation error to the low-rank approximation of certain trigonometric moment matrices, followed by a refined spectral analysis of their minimum eigenvalue. In the case of Gaussian mixing distributions, this result corrects a previous lower bound in [WV10b].
@@ 1 Introduction
Let $d(f,g)$ denote a loss function that measures the approximation error of $g$ by $f$. Concrete examples include $L_{p}$ distances or $f$-divergences [Csi75], the latter of which, including the total variation $\mathsf{TV}(f,g)$, squared Hellinger distance $H^{2}(f,g)$, the Kullback-Leibler divergence $\mathsf{KL}(f\|g)$, and the $\chi^{2}$-divergence $\chi^{2}(f\|g)$, are the focus of the present paper [^1] If $f$ and $g$ are densities for distribution $P$ and $Q$, respectively, we also write $d(P,Q)=d(f,g)$. (see Appendix A.1 for a quick review). The best approximation error of $f_{P}$ by an $m$-component mixture is defined as
@@
In information theory, the Gaussian convolution structure arises in the context of Gaussian channels [CT06], where the input and output distributions correspond to $P$ and $P*\phi$, respectively. The channel capacity determines the maximal rate at which information can be reliably transmitted, which, under the second moment constraint, is achieved by a Gaussian input distribution. In practice, the input may be constrained to be finitely valued due to modulation. To address this issue, [WV10b] studied the Gaussian channel capacity under input cardinality constraints (finite-constellation capacity), in particular, the rate of convergence to the Gaussian channel capacity when the cardinality grows. It turns out that this capacity gap is precisely characterized by ${\mathcal{E}}^{\star}$ under the KL divergence for the unconstrained capacity-achieving input distribution (Gaussian) – see (8).
@@
which offers a meaningful complexity measure for the class $\{f_{P}:P\in{\mathcal{P}}\}$ and is closely related to more classical complexity notions such as the metric entropy. In fact, most of the existing constructive bounds on the metric or bracketing entropy for general Gaussian mixtures are obtained by first finding a discrete approximation then quantizing the weights and atoms, and the resulting upper bounds increase with $m^{\star}$ [GvdV07a, GvdV01, Zha09, SG20]. Hence, tightened upper bounds for metric entropy immediately follow. See Section 5.1 for a detailed discussion.
@@
Next, we state our main results in terms of $m^{\star}$, which follow from ${\mathcal{E}}^{\star}$ in Sections 3 and 4.
@@ 1.1 Main results
Suppose $M\leq\epsilon^{-c_{1}}$ for some universal constant $10<c_{1}<$ 1. Then, for $\epsilon\in(0,1)$ and $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$,
@@
We provide some interpretations of Theorem 1. By definition, $m^{\star}$ increases with $M$ and decreases with $\epsilon$. In fact, (3) captures an “elbow-effect” depending on the relationship between $M$ and $\epsilon$. If the support of mixing distributions is not too wide, e.g., $M\lesssim\left(\log\frac{1}{\epsilon}\right)^{\frac{1}{2}-\delta}$ for a constant $\delta>0$, the complexity $m^{\star}$ has a slower growth with respect to $\epsilon$ as $\frac{\log\epsilon^{-1}}{\log\log\epsilon^{-1}}$. When $\sqrt{\log\frac{1}{\epsilon}}\lesssim M\lesssim\epsilon^{-c_{1}}$, the finite mixture needs to cover a wider range, and consequently, $m^{\star}$ grows faster as $M\sqrt{\log\frac{1}{\epsilon}}$. Finally, when $M\gtrsim\epsilon^{-c_{1}}$ for any $c_{1}\in(0,1)$, our Theorem 7 shows that we need at least $\Omega(M)$ components to encompass an extensive range.
Next, we turn to the families of distributions supported on the whole real line. The following notions are introduced to characterize the tail conditions. Given a function $\psi:[0,\infty)\mapsto[0,\infty)$ that is non-decreasing and satisfies that $\psi(0)=0$ and $\lim_{x\to\infty}\phi(x)=\infty$, define
@@
This concept is commonly referred to as the $\psi$-Orlicz norm, originally introduced in [KR61]. It also implies the following tail bound [KC22]:
@@
Suppose that $\beta\leq\epsilon^{-c_{1}}$ for some universal constant $10<c_{1}<$ 1. Then, for $\epsilon\in(0,1)$ and $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$,
@@
Theorem 2 shows that the complexity level grows in a logarithm rate, which equivalently implies an exponential convergence rate of the approximation error ${\mathcal{E}}^{\star}$. The tight result is applicable when $\alpha\in(0,2]$ and $\beta$ is a fixed constant, such as in the case of the subgaussian and subexponential family. Specifically, for the $\sigma^{2}$-subgaussian family ${\mathcal{P}}(2,\sigma)$ with $c_{0}\leq\sigma\leq\epsilon^{-c_{1}}$, we have
$m^{\star}(\epsilon,{\mathcal{P}}(2,\sigma),d)\asymp{\sigma\log\frac{1}{\epsilon}}.$ (7)
@@
We now briefly discuss the proof strategies for the main results. We prove the upper bound under the $\chi^{2}$-divergence and the lower bound under the $\mathsf{TV}$ distance. The results for $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$ quickly follow by the inequalities between $f$-divergences in Appendix A.1. For the upper bound, we extend the local moment matching argument in previous work [Zha09], which constructs a discrete approximation by matching the moments for the mixing distribution conditioned on each subinterval in a partition of the effective support of the mixture. This approach can be further generalized to distribution families characterized by various tail conditions. The matching lower bound is the major contribution of this paper, where we relate the best approximation error to the low-rank approximation of trigonometric moment matrices, and then conduct a refined analysis based on the classical spectrum theory. The application of orthogonal polynomials also plays a crucial role in our analysis.
Remark 1 (Gaussian location mixture with non-unitary variance).
Consider the problem of approximating $P*\phi_{\sigma}$ by an $m$-component mixture $P_{m}*\phi_{\sigma}$, where $\phi_{\sigma}$ denotes the density of $N(0,\sigma^{2})$. This problem appears in heteroscedastic settings of nonparametric density estimation [Jia20, SGS21]. By the scale invariance of the $f$-divergences, we have $d(P_{\sigma}*\phi,Q_{\sigma}*\phi)=d(P*\phi_{\sigma},Q*\phi_{\sigma})$ for $d\in\{\mathsf{TV},H,\mathsf{KL},\chi^{2}\}$, where $P_{\sigma}$ denotes the distribution of $X/\sigma$ for $X\sim P$. Consequently, the minimum number of components is equal to $m^{\star}(\epsilon,P_{\sigma},d)$.
@@ 1.2 Comparison with previous results
Below we give an overview of previous results. The upper bound for the compact support case is discussed in [GvdV01, GvdV07a, Zha09]. Among them, the strongest result [Zha09, Lemma 1] gives an upper bound of $m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{M},\mathsf{TV})\lesssim M\sqrt{\log\epsilon^{-1}}\vee\log\epsilon^{-1}$. Theorem 1 strengthens this result by bounding the $\chi^{2}$-divergence and establishing the optimal rate. Correspondingly, for general $\sigma$, it follows from Remark 1 that the $\frac{M}{\sigma}\sqrt{\log\epsilon^{-1}}\vee\log\epsilon^{-1}$ upper bound from [Jia20, Lemma 3] can be improved to $\frac{\log\epsilon^{-1}}{\log\left(1+\sigma\sqrt{\log\epsilon^{-1}}/{M}\right)}\vee 1.$ In addition, for Hellinger distance, [PW20] conjectures that $m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{1},H)\asymp\frac{\log\frac{1}{\epsilon}}{\log\log\frac{1}{\epsilon}}$, which is proved by our Theorem 1.
For the subgaussian case, [PW20, Lemma 7] gives a $\log\epsilon^{-1}$ upper bound for 1–subgaussian family, which is further extended to sub-Weibull families by Theorem 2. The specific problem of approximating $P=N(0,\sigma^{2})$ is studied in [WV10a, WV10b] in the context of the finite-constellation capacity. In fact,
${\mathcal{E}}^{\star}(m,N(0,\sigma^{2}),\mathsf{KL})=C-C_{m}$ (8)
where
$C=\max_{P_{X}:\mathbb{E}[X^{2}]\leq\sigma^{2}}I(X;X+Z)=\frac{1}{2}\log(1+\sigma^{2})$
is the Gaussian channel capacity with input $X$ and additive noise $Z\sim N(0,1)$, and
$C_{m}=\max_{P_{X}\in{\mathcal{P}}_{m}:\mathbb{E}[X^{2}]\leq\sigma^{2}}I(X;X+Z)$
is the capacity under the additional input cardinality constraint. While quantized Gaussian only achieves an error that is polynomial in $m$, an exponential upper bound is ${\mathcal{E}}^{\star}(m,N(0,\sigma^{2}),\mathsf{KL})\lesssim\sigma^{2}(\frac{\sigma^{2}}{1+\sigma^{2}})^{2m}$ is shown in [WV10b, Theorem 8] using the Gauss quadrature. As a corollary of our main result, (7) generalizes this result to subgaussian family with an improved exponent for large $\sigma$.
@@ 1.3 Related work
The problem of approximation by location mixtures is first addressed by the celebrated Tauberian theorem of Wiener [Wie32], which gives a general characterization of whether the translation family of a given function is dense in $L_{1}(\mathbb{R}^{d})$ or $L_{2}(\mathbb{R}^{d})$ in terms of its Fourier transform. Convergence rates have been studied over the past few decades, with a wide range of applications in approximation theory, machine learning, and information theory [Bar93, Fer97, CABR02, WV10b, WV10a]. For example, for the location and scale $m$-mixture class of sigmoidal functions, Barron [Bar93] obtained dimension-free convergence rate for approximating functions whose gradient has integrable Fourier transform, a fundamental result in the theory of neural networks. For Gaussian models, Wu and Verdú [WV10b] linked this problem to the Gaussian channel capacity under input cardinality constraint (cf. (8)). More recently, [NM19, NNCM20, NCNM22] showed the consistent approximation over various families with general location-scale mixtures. Another related problem is the Gaussian mixture reduction, which requires approximating a high-order Gaussian mixture with a low-order one. This problem broadly arises in applications including belief propagation and Bayesian filtering. Although there are many numerical algorithms by means of clustering, optimization, or the greedy algorithm, convergence rates and optimal approximators are still left to be discovered. See [ZZC23, SLP23, DMIO23] for some recent works and [CWPS11] for a review.
@@
In the statistics literature, understanding the complexity of a distribution class plays an important role in nonparametric density and functional estimation [vdVW96]. Information-theoretic risk bounds are obtained on the basis of metric entropy for a variety of loss functions (see [PW24, Chapter 32] for an overview). In addition, metric entropy of the Gaussian mixture family is crucial in the statistical analysis of the sieve and nonparametric maximum likelihood estimator (MLE) in mixture models as well as posterior concentration [SW94, WS95, GW00, GvdV01, GvdV07a, Zha09, SG20, Jia20, SGS21]. These results all rely on metric entropy of the Gaussian mixture class obtained via approximation by finite mixtures. In this vein, the quantity $m^{\star}(n^{-1/2},{\mathcal{P}},d)$ is referred to as the statistical degree of ${\mathcal{P}}$ [PW20], which represents the smallest $m$ so that any density in ${\mathcal{P}}$ can be made statistically indistinguishable (on the basis of $n$ observations) from some density in ${\mathcal{P}}_{m}$.
@@
A related quantity to finite mixture approximation, known as smoothed empirical empirical distribution, has been studied in recent literature [GGNWP20, CNW22, BJPR22]. Given a sample $(X_{1},\ldots,X_{m}){\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}}P$, we approximate $P*\phi$ by $P_{m}*\phi$, where $P_{m}=\frac{1}{m}\sum_{i=1}^{m}\delta_{X_{i}}$ is the empirical measure. For mixing distributions with bounded support or subgaussian tails, the convergence rate of this approximation is shown to be dimension-free and polynomial in $m$, measured by different Wasserstein distances and $f$-divergences (e.g., $O(m^{-1/2})$ under $\mathsf{TV}$ and $O(m^{-1})$ under $\chi^{2}$). Note that $P_{m}$ is a special $m$-atomic distribution. As noted in [WV10b, Sec. VII], the achieved approximation error is consistent with CLT-type of rates which are markedly slower than the optimal rates that are typically exponentially.

### v2

@@ Abstract
We consider the problem of approximating a general Gaussian location mixture by finite mixtures. The minimum order of finite mixtures that achieve a prescribed accuracy is determined within constant factors for the family of mixing distributions with compact support or appropriate assumptions on the tail probability including subgaussian and subexponential. While the upper bound is achieved using the technique of local moment matching, the lower bound is established by relating the best approximation error to the low-rank approximation of certain trigonometric moment matrices, followed by a refined spectral analysis of their minimum eigenvalue. In the case of Gaussian mixing distributions, this result corrects a previous lower bound in [WV10b].
@@ 1 Introduction
Let $d(f,g)$ denote a loss function that measures the approximation error of $g$ by $f$. Concrete examples include $L_{p}$ distances or $f$-divergences [Csi75], the latter of which, including the total variation $\mathsf{TV}(f,g)$, squared Hellinger distance $H^{2}(f,g)$, the Kullback-Leibler (KL) divergence $\mathsf{KL}(f\|g)$, and the $\chi^{2}$-divergence $\chi^{2}(f\|g)$, are the focus of the present paper [^1] If $f$ and $g$ are densities for distribution $P$ and $Q$, respectively, we also write $d(P,Q)=d(f,g)$. (see Appendix A.1 for a quick review). The best approximation error of $f_{P}$ by an $m$-component mixture is defined as
@@
In information theory, the Gaussian convolution structure arises in the context of Gaussian channels [CT06], where the input and output distributions correspond to $P$ and $P*\phi$, respectively. The channel capacity determines the maximal rate at which information can be reliably transmitted, which, under the second moment constraint, is defined as
$C\triangleq\max_{P_{X}:\mathbb{E}[X^{2}]\leq\sigma^{2}}I(X;X+Z),$
where the mutual information between two random variables $X$ and $Y$ is defined as $I(X;Y)\triangleq\mathbb{E}\log\frac{P_{XY}}{P_{X}P_{Y}}$. It is well-known that the capacity $C=\frac{1}{2}\log(1+\sigma^{2})$, and it is achieved by a Gaussian input distribution $N(0,\sigma^{2})$. While Gaussian inputs achieve the theoretical capacity, practical communication systems typically employ modulation schemes that restrict signals to finite and discrete constellations. This constraint introduces the problem of finite-constellation design, which aims to approach the theoretical capacity limit and improve system performance. Such design challenges frequently occur in diverse application domains, including wireless communication [WXD+18], quantum cryptography [KGW21, LUL19, GGDL19], and computational neuroscience [KL13].
To address this issue, [WV10b] studied the Gaussian channel capacity under input cardinality constraints (finite-constellation capacity), in particular, the rate of convergence to the Gaussian channel capacity when the cardinality grows. Define
$C_{m}\triangleq\max_{P_{X}\in{\mathcal{P}}_{m}:\mathbb{E}[X^{2}]\leq\sigma^{2}}I(X;X+Z)$
as the capacity subject to an input cardinality $m$. It turns out that the capacity gap is precisely characterized by ${\mathcal{E}}^{\star}$ under the KL divergence for the unconstrained capacity-achieving input distribution (Gaussian):
${\mathcal{E}}^{\star}(m,N(0,\sigma^{2}),\mathsf{KL})=C-C_{m}.$ (3)
Consequently, the rate of the capacity gap and the corresponding constellation design are direct applications of our main results. We construct a rate-optimal approximation using a refined moment-matching approach, which serves as a capacity-achieving constellation scheme for Gaussian channel inputs. Notably, we apply a local moment matching scheme that apply moment matching to the conditional distribution of Gaussian distribution on a sequence of intervals, which provably improves the global moment matching scheme in [WV10b]. (For further discussions and numerical comparison, see Section 5.1.) Furthermore, to understand the fundamental limit of the capacity gap, we establish a tight lower bound through a novel analytical framework in Section 4. Our analysis reduces the finite mixture approximation problem to a low-rank matrix approximation problem, which is resolved using spectral analysis. Notably, this lower bound corrects the previous result in [WV10b, Theorem 8].
@@
which offers a meaningful complexity measure for the class $\{f_{P}:P\in{\mathcal{P}}\}$ and is closely related to more classical complexity notions such as the metric entropy. In fact, most of the existing constructive bounds on the metric or bracketing entropy for general Gaussian mixtures are obtained by first finding a discrete approximation then quantizing the weights and atoms, and the resulting upper bounds increase with $m^{\star}$ [GvdV07, GvdV01, Zha09, SG20]. Hence, tightened upper bounds for metric entropy immediately follow. See Section 5.2 for a detailed discussion.
@@
In the following, we state our main results in terms of $m^{\star}$, which follow from bounds on ${\mathcal{E}}^{\star}$ in Sections 3 and 4.
@@ 1.1 Main results
We consider a loss function $d$ satisfying the following inequality. This assumption encompasses the $\mathsf{TV}^{2},H^{2},\mathsf{KL},\chi^{2}$ divergences as special cases by the inequalities between $f$-divergences (see Appendix A.1).
Assumption 1.
There exists absolute constants $c,c^{\prime}>0$ such that
$c\mathsf{TV}^{2}\left(P,Q\right)\leq d\left(P\|Q\right)\leq{c^{\prime}\chi^{2}\left(P\|Q\right)},\quad\forall P,Q.$
@@
Suppose $M\leq\epsilon^{-c_{1}}$ for some universal constant $0<c_{1}<\frac{1}{2}$. Then, for $\epsilon\in(0,\frac{1}{2}]$ and $d$ satisfying Assumption 1,
@@
We provide some interpretations of Theorem 1. By definition, $m^{\star}$ increases with $M$ and decreases with $\epsilon$. In fact, (4) captures an “elbow-effect” depending on the relationship between $M$ and $\epsilon$. If the support of mixing distributions is not too wide, e.g., $M\lesssim\left(\log\frac{1}{\epsilon}\right)^{\frac{1}{2}-\delta}$ for a constant $\delta>0$, the complexity $m^{\star}$ has a slower growth with respect to $\epsilon$ as $\frac{\log\epsilon^{-1}}{\log\log\epsilon^{-1}}$. When $\sqrt{\log\frac{1}{\epsilon}}\lesssim M\lesssim\epsilon^{-c_{1}}$, the finite mixture needs to cover a wider range, and consequently, $m^{\star}$ grows faster as $M\sqrt{\log\frac{1}{\epsilon}}$. Finally, when $M\gtrsim\epsilon^{-c_{1}}$ for any $c_{1}\in(0,\frac{1}{2})$, our Theorem 7 shows that we need at least $\Omega(M)$ components to encompass an extensive range.
Next, we turn to the families of distributions supported on the whole real line. The following notions are introduced to characterize the tail conditions. Given a function $\psi:[0,\infty)\mapsto[0,\infty)$ that is non-decreasing and satisfies that $\psi(0)=0$ and $\lim_{x\to\infty}\psi(x)=\infty$, define
@@
This concept is commonly referred to as the $\psi$-Orlicz norm, originally introduced in [MY61]. It also implies the following tail bound [KC22]:
@@
Suppose that $\beta\leq\epsilon^{-c_{1}}$ for some universal constant $0<c_{1}<\frac{1}{2}$. Then, for $\epsilon\in(0,\frac{1}{2}]$ and $d$ satisfying Assumption 1,
@@
Theorem 2 shows that the complexity level grows in a logarithm rate, which equivalently implies an exponential convergence rate of the approximation error ${\mathcal{E}}^{\star}$. The tight result is applicable when $\alpha\in(0,2]$ and $\beta$ is a fixed constant, such as in the case of the subgaussian and subexponential family. Specifically, for the $\sigma^{2}$-subgaussian family ${\mathcal{P}}_{2}(\sigma)$ with $c_{0}\leq\sigma\leq\epsilon^{-c_{1}}$, we have
$m^{\star}(\epsilon,{\mathcal{P}}_{2}(\sigma),d)\asymp{\sigma\log\frac{1}{\epsilon}}.$ (8)
In these theorems, the conditions $M,\beta\leq\epsilon^{-c_{1}}$ are required only for the lower bound of $m^{\star}$, and the upper bound continues to hold without this condition. For explicit details, see Theorems 3 and 7 for the corresponding upper and lower bounds.
@@
We now briefly discuss the proof strategies for the main results. We prove the upper bound under the $\chi^{2}$-divergence and the lower bound under the $\mathsf{TV}$ distance as $d$ satisfies Assumption 1. For the upper bound, we extend the local moment matching argument in previous work [Zha09], which constructs a discrete approximation by matching the moments for the mixing distribution conditioned on each subinterval in a partition of the effective support of the mixture. This approach can be further generalized to distribution families characterized by various tail conditions. The matching lower bound is the major contribution of this paper, where we relate the best approximation error to the low-rank approximation of trigonometric moment matrices, and then conduct a refined analysis based on the classical spectrum theory. The application of orthogonal polynomials also plays a crucial role in our analysis.
Remark 1 (Gaussian location mixture with general variance).
Consider the problem of approximating $P*\phi_{\sigma}$ by an $m$-component mixture $P_{m}*\phi_{\sigma}$, where $\phi_{\sigma}$ denotes the density of $N(0,\sigma^{2})$. This problem appears in heteroscedastic settings of nonparametric density estimation [Jia20, SGS24]. By the scale invariance of the $f$-divergences, we have $d(P_{\sigma}*\phi,Q_{\sigma}*\phi)=d(P*\phi_{\sigma},Q*\phi_{\sigma})$ for $d\in\{\mathsf{TV},H^{2},\mathsf{KL},\chi^{2}\}$, where $P_{\sigma}$ denotes the distribution of $X/\sigma$ for $X\sim P$. Consequently, the minimum number of components is equal to $m^{\star}(\epsilon,P_{\sigma},d)$.
Remark 2.
The loss function $d$ may not be symmetric, as in the case with $\mathsf{KL}$ and $\chi^{2}$ divergences. In general, the best approximation with respect to $d(f_{P_{m}},f_{P})$ and $d(f_{P},f_{P_{m}})$ are not equivalent. For instance, if $P=N(0,\sigma^{2})$ with $\sigma>1$, then $\chi^{2}(f_{P}\|f_{P_{m}})=\infty$ for any $m$ and any $m$-atomic $P_{m}$, but $\chi^{2}(f_{P_{m}}\|f_{P})$ can be made exponentially small.
In certain special cases, such as when $P$ is compactly supported or sub-Gaussian, the rate of $\mathsf{KL}(f_{P}\|f_{P_{m}})$ can be derived using the divergence comparison inequalities for Gaussian mixtures [JPW23]. These inequalities provide upper bounds on the $\mathsf{KL}$ divergence between Gaussian mixtures in terms of the symmetric squared Hellinger distance ($H^{2}$), where our main results can be applied.
@@ 1.2 Comparison with previous results
Below we give an overview of previous results. The upper bound for the compact support case is discussed in [GvdV01, GvdV07, Zha09]. Among them, the strongest result [Zha09, Lemma 1] gives an upper bound of $m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{M},\mathsf{TV})\lesssim\left(M\sqrt{\log{\frac{1}{\epsilon}}}\right)\vee\log{\frac{1}{\epsilon}}$. Theorem 1 strengthens this result by bounding the $\chi^{2}$-divergence and establishing the optimal rate. Correspondingly, for general $\sigma$, it follows from Remark 1 that the $\left(\frac{M}{\sigma}\sqrt{\log{\frac{1}{\epsilon}}}\right)\vee\log{\frac{1}{\epsilon}}$ upper bound from [Jia20, Lemma 3] can be improved to $\frac{\log{\frac{1}{\epsilon}}}{\log\left(1+\frac{\sigma}{M}\sqrt{\log{\frac{1}{\epsilon}}}\right)}\vee 1.$ In addition, for Hellinger distance, [PW20] conjectures that $m^{\star}(\epsilon,{\mathcal{P}}^{\textsf{Bdd}}_{1},H)\asymp\frac{\log\frac{1}{\epsilon}}{\log\log\frac{1}{\epsilon}}$, which is proved by our Theorem 1.
For the subgaussian case, [PW20, Lemma 7] gives a $\log{\frac{1}{\epsilon}}$ upper bound for 1–subgaussian family, which is further extended to sub-Weibull families by Theorem 2. The specific problem of approximating $P=N(0,\sigma^{2})$ is studied in [WV10a, WV10b] in the context of the finite-constellation capacity. While quantized Gaussian only achieves an error that is polynomial in $m$, an exponential upper bound is ${\mathcal{E}}^{\star}(m,N(0,\sigma^{2}),\mathsf{KL})\lesssim\sigma^{2}(\frac{\sigma^{2}}{1+\sigma^{2}})^{2m}$ is shown in [WV10b, Theorem 8] using the Gauss quadrature. As a corollary of our main result, (8) generalizes this result to subgaussian family with an improved exponent for large $\sigma$.
@@ 1.3 Related work
The problem of approximation by location mixtures is first addressed by the celebrated Tauberian theorem of Wiener [Wie32], which gives a general characterization of whether the translation family of a given function is dense in $L_{1}(\mathbb{R}^{d})$ or $L_{2}(\mathbb{R}^{d})$ in terms of its Fourier transform. Convergence rates have been studied over the past few decades, with a wide range of applications in approximation theory, machine learning, and information theory [Bar93, Fer97, CABR02, WV10b, WV10a]. For example, for the location and scale $m$-mixture class of sigmoidal functions, Barron [Bar93] obtained dimension-free convergence rate for approximating functions whose gradient has integrable Fourier transform, a fundamental result in the theory of neural networks. For Gaussian models, Wu and Verdú [WV10b] linked this problem to the Gaussian channel capacity under input cardinality constraint (cf. (3)). More recently, [NM19, NNCM20, NCNM22] showed the consistent approximation over various families with general location-scale mixtures. Another related problem is the Gaussian mixture reduction, which requires approximating a high-order Gaussian mixture with a low-order one. This problem broadly arises in applications including belief propagation and Bayesian filtering. Although there are many numerical algorithms by means of clustering, optimization, or the greedy algorithm, convergence rates and optimal approximators are still left to be discovered. See [ZZC24, SLP23, DMIO23] for some recent works and [CWPS11] for a review.
@@
In the statistics literature, understanding the complexity of a distribution class plays an important role in nonparametric density and functional estimation [vdVW96]. Information-theoretic risk bounds are obtained on the basis of metric entropy for a variety of loss functions (see [PW24, Chapter 32] for an overview). In addition, metric entropy of the Gaussian mixture family is crucial in the statistical analysis of the sieve and nonparametric maximum likelihood estimator (MLE) in mixture models as well as posterior concentration [SW94, WS95, GW00, GvdV01, GvdV07, Zha09, SG20, Jia20, SGS24]. These results all rely on metric entropy of the Gaussian mixture class obtained via approximation by finite mixtures. In this vein, the quantity $m^{\star}(n^{-1/2},{\mathcal{P}},d)$ is referred to as the statistical degree of ${\mathcal{P}}$ [PW20], which represents the smallest $m$ so that any density in ${\mathcal{P}}$ can be made statistically indistinguishable (on the basis of $n$ observations) from some density in ${\mathcal{P}}_{m}$.
@@
A related quantity to finite mixture approximation, known as smoothed empirical distribution, has been studied in recent literature [GGNWP20, CNW22, BJPR22]. Given a sample $(X_{1},\ldots,X_{m}){\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}}P$, we approximate $P*\phi$ by $P_{m}*\phi$, where $P_{m}=\frac{1}{m}\sum_{i=1}^{m}\delta_{X_{i}}$ is the empirical measure. For mixing distributions with bounded support or subgaussian tails, the convergence rate of this approximation is shown to be dimension-free and polynomial in $m$, measured by different Wasserstein distances and $f$-divergences (e.g., $O(m^{-1/2})$ under $\mathsf{TV}$ and $O(m^{-1})$ under $\chi^{2}$). Note that $P_{m}$ is a special $m$-atomic distribution. As noted in [WV10b, Sec. VII], the achieved approximation error is consistent with CLT-type of rates which are markedly slower than the optimal rates that are typically exponentially.